### v2

# StableMaterials: Enhancing Diversity in Material Generation via Semi-Supervised Learning

Abstract.
We introduce StableMaterials, a novel approach for generating photorealistic physical-based rendering (PBR) materials that integrate semi-supervised learning with Latent Diffusion Models (LDMs). Our method employs adversarial training to distill knowledge from existing large-scale image generation models, minimizing the reliance on annotated data and enhancing the diversity in generation. This distillation approach aligns the distribution of the generated materials with that of image textures from an SDXL model, enabling the generation of novel materials that are not present in the initial training dataset. Furthermore, we employ a diffusion-based refiner model to improve the visual quality of the samples and achieve high-resolution generation. Finally, we distill a latent consistency model for fast generation in just four steps and propose a new tileability technique that removes visual artifacts typically associated with fewer diffusion steps.
We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks, and show the advantages of our approach. Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond.
StableMaterials is publicly available at https://gvecchio.com/stablematerials.
material appearance, generative models
journal: TOGccs: Computing methodologies Computer graphics
Figure 1. We present StableMaterials, a diffusion–based model for materials generation through text or image prompting. Our approach enables high-resolution, tileable material maps, inferring both diffuse (Basecolor) and specular (Roughness, Metallic) properties, as well as the material mesostructure (Height, Normal).
[…]
Teaser Figure
1. Introduction
Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise. To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2023], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2023, 2024]. While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
Despite recent efforts to create large-scale materials datasets, such as Deschaintre et al. [2018], OpenSVBRDF [Ma et al., 2023], and MatSynth [Vecchio and Deschaintre, 2024], these datasets are limited in diversity [Zhou et al., 2023], not capturing the vast range observed in large-scale image datasets such as LAION [Schuhmann et al., 2022]. These limitations can constrain the capabilities of learning-based approaches, potentially creating gaps in their generative capabilities and affecting realism and diversity.
Fine-tuning has become a common practice to reduce these gaps in training data by leveraging existing knowledge from large-scale pretrained models. Techniques like Low-Rank Adaptation (LoRA) [Hu et al., 2021] effectively fine-tune models while preventing catastrophic forgetting. Methods such as Diff-Instruct [Luo et al., 2024], on the other hand, employ distillation strategies to transfer knowledge from pretrained models. However, while fine-tuning or distillation within the same domain are straightforward, they pose significant challenges across different domains (e.g., image to material).
To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [Podell et al., 2023] model. In particular, we use a pretrained SDXL to generate unannotated texture samples from text prompts. However, since StableMaterials is trained to produce SVBRDF maps, we cannot perform direct supervision using the generated textures. To include these textures in the training of our methods, we learn a common latent representation between textures and materials; then, we complement the traditional supervised loss, with an unsupervised adversarial loss, forcing the model to also generate realistic maps for unannotated samples and close the gap between the two data distributions. In addition, drawing inspiration from the SDXL [Podell et al., 2023] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation. We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [Meng et al., 2021] and patched diffusion. This approach allows to achieve high resolution while constraining the patched generation, ensuring consistency and memory efficiency. Subsequently, we distill a latent consistency model [Song et al., 2023] that allows fast generation by reducing the number of inference steps to four steps per stage. However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2023] to achieve tileability. To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
We qualitatively evaluate our method and compare it with previous work, demonstrating the benefit of our semi-supervised training approach. In summary, we introduce StableMaterials a novel solution combining supervised and adversarial training to generate highly realistic material in scenarios where annotated data are scarce. The contributions of this work are as follows:
  • StableMaterials, a new diffusion–based model for PBR material generation, leveraging a semi-supervised learning approach to incorporate unannotated data during training.
  • A novel distillation technique to bridge the gap with large-scale models by including unannotated data in the training.
  • A novel “features rolling” approach to tileability, minimizing the visual artifacts produced by fewer diffusion steps.
  • State-of-the-art performance in PBR materials generation.
2. Related Work
Figure 2. Architecture of StableMaterials. The base model generates a low resolution materials of size 512x512. This generation is then upscaled and refined using SDEdit [Meng et al., 2021] by the refiner model using a patched approach to limit memory requirements.
[…]
Method Figure
Materials Generation.
Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2023]. Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al. [2020] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [Guo et al., 2020], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [Karras et al., 2020a]; Hu et al. [2022] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [Zhou et al., 2022], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training. Recent appraoches have focused on leveraging the generative capabilities of diffusion models for materials generation. In particular, Vecchio et al. [2023] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image. MatFuse [Vecchio et al., 2024] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting. MaterialPalette [Lopes et al., 2024] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [Hu et al., 2021] for each picture. Substance 3D Sampler [Adobe, 2024] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning. However, these methods often lack diversity, struggling with complex material representation, or depend on image generation models, requiring additional steps to estimate the material parameters. StableMaterials overcomes these limitations by including a wider variety of unannotated material samples via a semi-supervised training and improves on inference time by distilling a latent consistency model.
Generative models.
Image generation is a long-standing challenge in computer vision, primarily due to the high dimensionality of images and complex data distributions. Generative Adversarial Networks (GAN) [Goodfellow et al., 2014] enabled the creation of high-quality images [Karras et al., 2017; Brock et al., 2018; Karras et al., 2020b], yet they suffer from unstable training [Arjovsky et al., 2017; Gulrajani et al., 2017; Mescheder, 2018], leading to mode collapse behaviour.
Diffusion Models (DMs) [Sohl-Dickstein et al., 2015; Ho et al., 2020], particularly the more efficient Latent Diffusion (LDM) architecture [Rombach et al., 2022], have emerged as an alternative to GANs, achieving state-of-the-art results in image generation tasks [Dhariwal and Nichol, 2021], besides showing a more stable training. Following the success of LDMs, research has focused on improving the generation quality [Podell et al., 2023] and reducing the number of inference steps to speed up the generation process [Song et al., 2023; Luo et al., 2023; Sauer et al., 2023b, 2024] through model distillation. Furthermore, due to the proliferation of large-scale pretrained models, several approaches were proposed to reuse their knowledge [Hu et al., 2021; Luo et al., 2024; Ruiz et al., 2023]. However, these approaches focus on fine-tuning within the same image domain, restricting their applicability to non-image domains.
3. Method
StableMaterials’ architecture, shown in Fig. 2, is based on MatFuse [Vecchio et al., 2024], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality pixel-level reflectance properties for arbitrary materials. We modify the original MatFuse architecture, which employed a multi-encoder VAE to learn a disentangled latent representation of the maps, by fine-tuning a resource-efficient single-encoder compression model with the same latent properties. Additionally, we introduce a semi-supervised training strategy to distill knowledge from a large-scale SDXL model and improve generation diversity. We employ latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast tileble generation, and achieve high resolution via a refiner model.
3.1. Material Representation
StableMaterials generates materials in the form of SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the material mesostructure. In particular, we generate base color, normal, height, roughness and metalness properties, with the roughness determining the BRDF specular lobe width and the metalness defining which area of the material represents a conductor.
3.2. Material Generation
The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a compact latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
Map Compression
We first train a multi-encoder VAE, consisting of a set of encoders =$\mathcal{E}=\left\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\right\}$ and a decoder $\mathcal{D}$, to encode $N$ material maps $M=\left\{\textbf{M}_{1},\textbf{M}_{2},\dots,\textbf{M}_{N}\right\}$ into their latent representation $z=concat\left(z_{1},z_{2},\dots,z_{N}\right)$, with $z_{i}=\mathcal{E}_{i}(\textbf{M}_{i})$ being the encoding of the $i^{th}$ map using the $i^{th}$ encoder $\mathcal{E}_{i}$.
The latent tensor $z$ is then decoded to reconstruct the input maps $\hat{M}=\mathcal{D}(z)$. Following [Vecchio et al., 2024], the VAE is trained using pixel-space $L_{2}$ loss, perceptual LPIPS loss [Zhang et al., 2021], a patch-based adversarial objective [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018]. We regularize the VAE latent space using a Kullback-Leibler divergence penalty [Kingma and Welling, 2013; Rezende et al., 2014].
Following the training of the multiencoder VAE, we fine-tune a resource-efficient single encoder model. In particular, we freeze the decoder and train only the maps encoder $\mathcal{E}^{\prime}$ to compress the concatenated material maps to the same latent representation $z$. This encoder maintains the original’s disentangled latent representation and compression efficiency with fewer parameters. We compare the two models in Sec. 4.4.1.
Finally, we learn a common latent representation between texture images and materials, allowing us to incorporate these unannotated textures into the training process. To do so, we fine-tune an autoencoder model to compress a single texture (i.e.: the basecolor map) into $z$. Similarly to the single-encoder model, we keep the decoder frozen and train the encoder only. This network is used for semi-supervised training, allowing us to inject unannotated textures generated using SDXL in the training. We detail the use of this model in Sec. 3.3.
Diffusion Model
We train a diffusion model on a compressed latent representation $z$ of the material. This model, based on the architecture of Rombach et al. [2022], uses a time-conditional U-Net [Ronneberger et al., 2015] to denoise the latent vectors $z$. During training, we generate noised latent vectors using a deterministic forward diffusion process $q\left(z_{t}|z_{t-1}\right)$, as defined in Ho et al. [2020], transforming them into an isotropic Gaussian distribution. The diffusion network $\epsilon_{\theta}$ learns the backward diffusion $q(z_{t-1}|z_{t})$ to denoise and reconstruct the original latent vector. The model training is described in Sec. 3.3
Conditioning
StableMaterials enables control over the generation via high-level text or image prompts, descriptive of the global appearance of the material. Conditions are encoded into a one-dimensional feature vector using a pretrained CLIP [Radford et al., 2021] model. During training, we alternate between images and text prompts to enhance the robustness of the model. In particular, we use an ambient-lit rendering of the material as the image condition and the caption as text prompt. For materials without a caption, we use descriptive tags to build a text prompt. For each training batch, we select between image and text conditioning modes by randomly dropping one modality. A drop probability of 75% for text and 25% for image empirically proved to be the most effective.
3.3. Semi-Supervised Adversarial Distillation
To reduce the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model. However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties representation) during training. In particular, we complement the supervised loss with an unsupervised adversarial loss through a discriminator on the latent space ($LD$). This adversarial loss forces the generator to synthesize materials that present features similar to ground-truth ones, while the discriminator learns to distinguish between the two data sources. The generator is trained using samples from the material dataset as well as unannotated textures, while the discriminator is trained using only annotated samples.
Our generator is optimized using a combination of supervised and adversarial losses. The supervised loss, $\mathcal{L}_{\text{sup}}$, is defined as:
(1) $\mathcal{L}_{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\left[\left\|\epsilon_{t},\epsilon_{\theta}\left(z_{t,\text{mat}},t\right)\right\|^{2}\right]+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\left[\left\|\epsilon_{t},\epsilon_{\theta}\left(z_{t,\text{tex}},t\right)\right\|^{2}\right]$
comparing the model-estimated noise $\epsilon\theta(z_{t},t)$ with the actual noise $\epsilon_{t}$ introduced during the forward diffusion process [Ho et al., 2020]. Here, $z_{t,mat}$ and $z_{t,tex}$ are the noisy latent at timestep $t$ for the materials and textures samples respectively. The loss computed on unannotated texture samples is weighted using the hyperparameter $\alpha$, set to $0.15$ in our experiments.
The adversarial objective $\mathcal{L}_{\text{adv}}$ push the generator to close the distribution gap between materials and textures, producing outputs that are perceived as materials by the discriminator. It is computed on the denoised latents at timestep $t-1$, and is defined as:
(2) $\mathcal{L}_{\text{adv}}=-\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z})}\left[LD(z_{t-1})\right]$
with $z_{t-1}=\text{concat}\left(z_{t-1,mat},z_{t-1,tex}\right)$ the concatenated denoised latents and $LD(\cdot)$ the output of the latent discriminator.
Following Sauer et al. [2021, 2023a], $LD$ is trained using a hinge loss function [Lim and Ye, 2017] computed between the noised latent embeddings of the materials $z_{t-1,\text{real}}$ encoded with the VAE, and $z_{t-1,\text{fake}}$ the denoised latents from the generator.
$LD$ shares the same architecture as the conditional diffusion U-Net encoder and is initialized with the weights of the diffusion U-Net. Functioning in a time-conditional manner, it incorporates the diffusion timestep and the CLIP embeddings from either text or image prompts, ensuring that evaluations are relevant to the desired material appearance.
Figure 3. Semi-Supervised training. Both the SDXL model and StableMaterials are prompted to generate the same material. The supervised $\mathcal{L}_{2}$ loss is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
[…]
Method Figure
3.4. Fast High-Resolution Generation
Few steps generation
To enhance generation efficiency, we finetune a Latent Consistency Model [Luo et al., 2023] (LCM). This model streamlines the denoising process by implementing a one-stage guided distillation of an augmented probability flow ordinary differential equation (PF-ODE). The distillation leverages a consistency function $f_{\theta}:(z_{t},c,t)\mapsto z_{0}$, which directly predicts the PF-ODE solution at $t=0$. Compared to two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the augmented PF-ODE. Additionally, LCM adopts a skipping step strategy, ensuring consistency not between consecutive steps, but between steps separated by $k$, thus $t_{n+k}\rightarrow t_{n}$, circumventing alignment issues in traditional models. The use of consistency models enables few-steps generation, resulting in fast material synthesis as detailed in Sec. 4.2.
Features rolling
Recent material generation methods achieve tileability via noise rolling [Vecchio et al., 2023], leveraging the iterative nature of diffusion models. However, this approach becomes ineffective with fewer generation steps, being unable to fully remove seams. Circular padding [Zhou et al., 2022], on the other hand, despite being independent of diffusion steps, is not suitable for patched generation as it causes self-tileability in patches.
To address these issues, we propose shifting tensor rolling from the diffusion step to the U-Net architecture by directly rolling the feature maps. This method, named features rolling, integrates rolling within U-Net blocks, reducing the need for multiple diffusion steps. By randomly translating and then reversing feature maps within each convolution and attention layer in the U-Net, this approach preserves edge continuity and minimizes visual artifacts typically associated with fewer diffusion steps. To better preserve structure in highly structured materials, we enable features rolling after the first diffusion step. We compare features rolling with other approaches to tileability in the Supplemental Materials.
Latent Upscaling & Refinement
Features rolling in combination with patch diffusion allows efficient high-resolution generation; however, it results in inconsistent generations between patches, due to the loss of cross-patch dependencies.
To overcome this limitation, we use a two-stage pipeline, similar to the one introduced in SDXL [Podell et al., 2023], where the output of the base model is further refined by a high-resolution expert using SDEdit [Meng et al., 2021]. This approach allows us to use patched diffusion in the refinement process while keeping the overall content consistent with the original generation. In particular, we first generate at the model’s base 512x512 resolution, and then refine it at the desired resolution. We apply the SDEdit refinement with a strength of $0.5$, which we find to be the best compromise between new high-frequency details cretion and consistency with the previous step generation. To better learn the high-frequency features of the materials, the refiner model is trained on 512x512 crops extracted from the 4K materials at their full resolution without downsampling. We show the quality improvement of this approach in the ablation study, Sec. 4.4.3, and provide additional samples in Supplemental Materials.
4. Implementation & Results
In this section, we first introduce the datasets used in our work, we then assess the generation capabilities of StableMaterials and compare it with recent state-of-the-art methods. Finally, we evaluate our design choices in the ablation studies.
4.1. Dataset
We train our model on the combined MatSynth [Vecchio and Deschaintre, 2024] and Deschaintre et al. [2018] datasets, for a total of $6,198$ unique PBR materials, using the original training/test splits. Our dataset includes 5 material maps (Basecolor, Normal, Height, Roughness, Metallic) and their renderings under different environmental illuminations. We complement the dataset with $4,000$ texture-text pairs from SDXL [Podell et al., 2023], using $200$ prompts. We query a ChatGPT [OpenAI, 2024] model to suggest relevant material prompts. The full list of prompts used, as well as samples generated from SDXL, are provided in the Supplemental Material.
4.2. Technical details
We train all models on a single NVIDIA RTX4090 GPU with 24GB of VRAM, employing gradient accumulation to achieve a larger batch.
Autoencoder
The compression model is trained with mini-batch gradient descent, using the Adam optimizer [Kingma and Ba, 2014] and a batch size of $8$. We train the model for $1,000,000$ iterations with a learning rate of $4.5\cdot 10^{-4}$ and enable the $\mathcal{L}_{\text{adv}}$ after $300,000$ iterations as in Esser et al. [2021]. We first train a multi-encoder VAE [Vecchio et al., 2024]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab. 1). We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
Latent Diffusion model
The diffusion model is trained supervisedly for $400,000$ iterations with a batch size of $16$ using an AdamW [Loshchilov and Hutter, 2017] optimizer, with a learning rate of $3.2\cdot 10^{-5}$. The model is fine-tuned semi-supervisedly for $200,000$ iterations with a batch size of $8$. We fine-tune the refiner model for $50,000$ iterations using a batch size of $16$ and a learning rate of $2\cdot 10^{-7}$. Both models use the original OpenAI U-Net architecture with 18 input and output channels.
Prompt Basecolor Normal Height Rough. Metallic Render Prompt Basecolor Normal Height Rough. Metallic Render
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4. Image-prompting. StableMaterials accurately captures the visual appearance of the input image, producing realistic materials for both in-domain (on the left) and out-domain (on the right) prompts. The render highlights the model’s ability to handle diverse and complex surfaces.
Basecolor Normal Height Rough. Metallic Render Basecolor Normal Height Rough. Metallic Render
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
‘Old rusty metal bars.’ ‘Fish scales reflecting a multitude of colors .’
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
‘Old wooden parquet floor.’ ‘Batik fabric with Indonesian patterns.’
Figure 5. Text-prompting. StableMaterials closely follow the input prompt, producing realistic materials for both in-domain (on the left) and out-domain (on the right) samples. The render highlights the model’s ability to generate accurate properties for different types of materials.
Prompt MatFuse MatGen Material Palette Stable Materials Prompt MatFuse MatGen Material Palette Stable Materials
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6. Comparison for image-prompting. We compare StableMaterials with MatFuse, MatGen, and MaterialPalette in image-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model. StableMaterials improves over previous methods quality and ability to captures the visual appearance of the input image. Full set of maps in Supplemental Materials.
MatFuse MatGen Substance 3D Sampler Stable Materials MatFuse MatGen Substance 3D Sampler Stable Materials
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
‘Cobblestone walkway covered in snow. Stones in different sizes and colors.’ ‘Italian velvet fabric with Renaissance art inspired prints.’
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
‘Volcanic lava.’ ‘Coral pavona with a bumpy and porous texture.’
Figure 7. Comparison for text prompting. We compare StableMaterials with MatFuse, MatGen, and Substance 3D Sampler on text-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model. StableMaterials better follows the input prompt and successfully models out-domain materials. Visual quality is on par or better than models trained on larger datasets. Full set of maps in Supplemental Materials.
Latent Consistency model
The latent consistency model [Luo et al., 2023] is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$. We use a linear schedule for $\beta$ and denoise using the LCM (Latent Consistency Models) sampling schedule at inference time with $T=4$ steps.
Inference
We assess execution speed and memory usage. We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel. Both base and refiner models run at half-precision. Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM. For comparison, the LDM with a DDIM sampler and 50 generation steps, followed by 25 refinement steps, takes 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
4.3. Results and comparison
For all results, we show generated material maps, as well as the rendering under ambient lighting. We include a wider variety results in the Supplemental Material, as well as material editing samples and a CLIP-based Nearest-Neighbour search in the training database.
Generation results
We evaluate our model generation capabilities for both image (Fig. 4 and Fig. 12) and text (Fig. 5 and Fig. 13) conditioning. The results show the ability of StableMaterials to closely follow the input prompt, producing diverse realistic materials. For each set of results, we show two in-domain generations, that is materials categories that are present in our annotated training dataset, and two out-domain samples, materials from the unannotated data, confirming the ability of StableMaterials to generate new high-quality materials from both data distributions.
Comparison
We compare StableMaterials against MatFuse [Vecchio et al., 2024] trained on MatSynth [Vecchio and Deschaintre, 2024], MatGen [^1] Generation samples for MatGen were kindly provided by the authors.  [Vecchio et al., 2023], Material Palette [Lopes et al., 2024], and the text-to-texture-to-material pipeline in Adobe Substance 3D Sampler [Adobe, 2024]. Figures 6 and 7 show the generation results for image and text conditioning, respectively. It should be noted that MatFuse, Material Palette, and StableMaterials are trained on public datasets, while both MatGen and Sampler use private data. Furthermore, MatFuse, MatGen, and StableMaterials directly produce PBR maps, while MaterialPalette and Sampler employ a two-step process: first, generating a texture, then using an SVBRDF estimator to predict the material maps. This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exarcebated by the non-material-specific training sometimes generating natural images (with perspective) instead of surfaces. Moreover, MaterialPalette fine-tunes a LoRA for each prompt, introducing a time and computation overhead. In contrast, StableMaterials directly creates material maps by having learned material features, leading to more precise and unbiased results. MatFuse is constrained by its $256\times 256$ resolution, resulting in blurry, undetailed generations, and struggles with complex materials. MatGen, on the other hand, generally synthesize high-quality materials, but presents oversharpening artifacts and struggles in following more complex promps. We only show the generation renderings in Figures 6 and 7, with the PBR maps and additional samples included in the Supplemental Materials.
4.4. Ablation Study
We evaluate our different design choices by comparing the performance of our model against the baseline solutions. We provide high-resolution ablation results in the Supplemental Materials.
4.4.1. VAE Architecture
[…]
Basecolor Normal Height Rough. Metallic
$-\mathcal{E}$ (271M par.) 0.030 0.035 0.030 0.032 0.016
$-\mathcal{E}$ (101M par.) 0.028 0.037 0.030 0.032 0.015
We report the reconstruction performance, in terms of RMSE, for both the multi-encoder and single-encoder models in Tab. 1. Results show that the adopted transfer-learning strategy allows us to obtain a smaller network while achieving performances comparable to those of the larger multiencoder VAE and retaining the same dientangled latent space.
4.4.2. Training strategies
We evaluate the semi-supervised training on model’s generation capabilities in Fig. 8. The base model (w/o semi-supervision) generates high-quality in-domain material but, as expected, struggles with out-domain materials absent from the original training data. The semi-supervised approach, in contrast, enhances diversity and overall quality, allowing the generation of materials beyond the initial training scope.
In-domain samples Out-domain samples
W/o Semi-supervision W/ Semi-supervision W/o Semi-supervision W/ Semi-supervision
Refer to caption Refer to caption Refer to caption Refer to caption
‘Terracotta brick wall with white grout.’ ‘Kilim Rugs carpets with complex geometric patterns.’
Figure 8. Ablation study of the training stategies. The model effectively generates high-quality material with image or text prompts but struggles with unrepresented materials. Semi-supervised learning improves generation quality and diversity, including new materials.
Base Refined Base Refined
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9. Refinement ablation results. Generation crops: the use of a diffusion refiner significantly enhances generation quality and sharpness.
4.4.3. High-resolution generation
We evaluate our method with and without the refiner in Fig. 9. The generation looks sharper, less noisy, and more detailed. We demonstrate the effectiveness of our two-stage approach to high-resolution generation in Fig. 10, comparing it to patched diffusion, complemented by feature rolling. Patched diffusion leads to increased scale (pattern replication) rather than enhanced resolution (more pixel density) and causes inconsistencies like scale variations between patches. In contrast, our approach generates consistent and better quality high-res materials by refining the generation at the model’s native resolution of $512\times 512$ pixels via a second diffusion step at the target resolution.
Patch diffusion Two-stages pipeline
Refer to caption Refer to caption
‘Tortoise shell.’
Figure 10. Ablation of the approaches to high-resolution. (1) Patched diffusion produces more scale rather than resolution and presents inconsistency in scale. (2) Our two-stage approach generates at the model’s native resolution, and refines the material.
5. Limitations and Future Work
StableMaterials presents some limitations as shown in Fig. 11. First, it struggles to handle natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures. Introducing more variety in training prompts could help mitigate the problem. Additionally, it can occasionally generate incorrect reflectance properties (e.g.: material misclassified as metal) for material classes that are only present in the unannotated dataset. The use of text prompts describing surface properties, at training time, could mitigate the problem.
(1) (2) (3)
Refer to caption Refer to caption Refer to caption
‘Squared tiles enclosed by rectangular tiles.’ ‘Yukata fabric with patterns of dragons.’ ‘Shoji Screens made of paper and wood.’
Figure 11. Limitations. Left to right: (1) Struggles with complex prompts describing spatial relations. (2) Unable to represent complex figures or patterns. (3) Can hallucinate reflectance properties.
6. Conclusion
We present StableMaterials, a novel diffusion model for fast generation of tileable high-resolution materials. Our model leverages semi-supervised training and knowledge distillation from large-scale pre-trained models to overcome the limitations in training data and improve material realism and variety. This approach significantly reduces reliance on annotated data, thereby enriching the diversity and realism of generated PBR materials.
In conclusion, StableMaterials presents a robust solution to material generation challenges and we believe it can serve as a blueprint for future research to harness the potential of large-scale pretrained models and unsupervised data in computer graphics and beyond.

### v3

@@
 Finally, we distill a latent consistency model for fast generation in just four steps and propose a new tileability technique that removes visual artifacts typically associated with fewer diffusion steps.
-We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks, and show the advantages of our approach.
+We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks.
 Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond.
-StableMaterials is publicly available at https://gvecchio.com/stablematerials.
+StableMaterials will be made publicly available.
 material appearance, generative models
-journal: TOGccs: Computing methodologies Computer graphics
+journal: TOGccs: Computing methodologies Appearance and texture representations
 Figure 1.
@@
 Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise.
-To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2023], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2023, 2024].
+To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2024a, b].
 While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
@@
 Subsequently, we distill a latent consistency model [Song et al., 2023] that allows fast generation by reducing the number of inference steps to four steps per stage.
-However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2023] to achieve tileability.
+However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2024a] to achieve tileability.
 To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
-We qualitatively evaluate our method and compare it with previous work, demonstrating the benefit of our semi-supervised training approach.
+We evaluate our method qualitatively and quantitalively, and compare it with previous work, demonstrating the benefit of the semi-supervised training.
 In summary, we introduce StableMaterials a novel solution combining supervised and adversarial training to generate highly realistic material in scenarios where annotated data are scarce.
@@
 Materials Generation.
-Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2023].
+Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a].
 Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al.
@@
 In particular, Vecchio et al.
-[2023] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image.
-MatFuse [Vecchio et al., 2024] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting.
+[2024a] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image.
+MatFuse [Vecchio et al., 2024b] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting.
 MaterialPalette [Lopes et al., 2024] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [Hu et al., 2021] for each picture.
@@
 3. Method
-StableMaterials’ architecture, shown in Fig.
-2, is based on MatFuse [Vecchio et al., 2024], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality pixel-level reflectance properties for arbitrary materials.
-We modify the original MatFuse architecture, which employed a multi-encoder VAE to learn a disentangled latent representation of the maps, by fine-tuning a resource-efficient single-encoder compression model with the same latent properties.
-Additionally, we introduce a semi-supervised training strategy to distill knowledge from a large-scale SDXL model and improve generation diversity.
-We employ latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast tileble generation, and achieve high resolution via a refiner model.
+StableMaterials builds on MatFuse [Vecchio et al., 2024b], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality, pixel-level reflectance properties for arbitrary materials (Fig.
+2).
+We replace MatFuse’s multi-encoder VAE architecture (used to learn a disentangled latent representation of material maps) with a more resource-efficient single-encoder model, fine-tuned to preserve the same latent properties.
+In addition, we introduce a semi-supervised training strategy that distills knowledge from a large-scale SDXL model to increase generation diversity.
+Our method leverages latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast, tileable generation.
+A dedicated refiner model enables high-resolution outputs while preserving global consistency.
 3.1. Material Representation
-StableMaterials generates materials in the form of SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the material mesostructure.
-In particular, we generate base color, normal, height, roughness and metalness properties, with the roughness determining the BRDF specular lobe width and the metalness defining which area of the material represents a conductor.
+StableMaterials generates SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the mesostructure of the material.
+Specifically, the generated maps include base color, normal, height, roughness, and metalness, where roughness specifies the specular lobe width and metalness indicates conductor regions.
 3.2. Material Generation
-The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a compact latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
+The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
 Map Compression
-We first train a multi-encoder VAE, consisting of a set of encoders =$\mathcal{E}=\left\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\right\}$ and a decoder $\mathcal{D}$, to encode $N$ material maps $M=\left\{\textbf{M}_{1},\textbf{M}_{2},\dots,\textbf{M}_{N}\right\}$ into their latent representation $z=concat\left(z_{1},z_{2},\dots,z_{N}\right)$, with $z_{i}=\mathcal{E}_{i}(\textbf{M}_{i})$ being the encoding of the $i^{th}$ map using the $i^{th}$ encoder $\mathcal{E}_{i}$.
-The latent tensor $z$ is then decoded to reconstruct the input maps $\hat{M}=\mathcal{D}(z)$.
-Following [Vecchio et al., 2024], the VAE is trained using pixel-space $L_{2}$ loss, perceptual LPIPS loss [Zhang et al., 2021], a patch-based adversarial objective [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018].
-We regularize the VAE latent space using a Kullback-Leibler divergence penalty [Kingma and Welling, 2013; Rezende et al., 2014].
-Following the training of the multiencoder VAE, we fine-tune a resource-efficient single encoder model.
-In particular, we freeze the decoder and train only the maps encoder $\mathcal{E}^{\prime}$ to compress the concatenated material maps to the same latent representation $z$.
-This encoder maintains the original’s disentangled latent representation and compression efficiency with fewer parameters.
-We compare the two models in Sec.
+We first train a multi-encoder VAE with encoders $\mathcal{E}={\mathcal{E}_{1},\dots,\mathcal{E}_{N}}$ and a decoder $\mathcal{D}$.
+Each encoder $\mathcal{E}_{i}$ encodes the $i^{th}$ map $\textbf{M}_{i}$ into a latent vector $z_{i}$, and the concatenated tensor $z=concat(z_{1},\dots,z_{N})$ is decoded to reconstruct the maps $\hat{M}=\mathcal{D}(z)$.
+Following [Vecchio et al., 2024b], training uses pixel-space $L_{2}$ loss, perceptual LPIPS [Zhang et al., 2021] loss, a patch-based adversarial loss [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018], with a Kullback-Leibler penalty [Kingma and Welling, 2013; Rezende et al., 2014] to regularize the latent space.
+Afterward, we fine-tune a single-encoder model.
+We freeze the original decoder and train only the encoder $\mathcal{E}^{\prime}$ to compress concatenated material maps into the same latent tensor $z$.
+This preserves the disentangled latent representation and maintains compression efficiency with fewer parameters, as shown in Sec.
 4.4.1.
-Finally, we learn a common latent representation between texture images and materials, allowing us to incorporate these unannotated textures into the training process.
-To do so, we fine-tune an autoencoder model to compress a single texture (i.e.: the basecolor map) into $z$.
-Similarly to the single-encoder model, we keep the decoder frozen and train the encoder only.
-This network is used for semi-supervised training, allowing us to inject unannotated textures generated using SDXL in the training.
-We detail the use of this model in Sec.
+Lastly, we create a shared latent space for texture and materials.
+This allows us to have a common space that we can use to train our diffusion model.
+Specifically, we fine-tune an autoencoder that compresses a single texture (e.g., the base color) into $z$, again keeping the decoder frozen.
+This additional network is used for the semi-supervised training of the diffusion model, as described in Sec.
 3.3.
 Diffusion Model
-We train a diffusion model on a compressed latent representation $z$ of the material.
-This model, based on the architecture of Rombach et al.
+We train a diffusion model on the compressed latent representation $z$ of the material.
+This model, based on Rombach et al.
 [2022], uses a time-conditional U-Net [Ronneberger et al., 2015] to denoise the latent vectors $z$.
@@
 Conditioning
-StableMaterials enables control over the generation via high-level text or image prompts, descriptive of the global appearance of the material.
-Conditions are encoded into a one-dimensional feature vector using a pretrained CLIP [Radford et al., 2021] model.
-During training, we alternate between images and text prompts to enhance the robustness of the model.
-In particular, we use an ambient-lit rendering of the material as the image condition and the caption as text prompt.
-For materials without a caption, we use descriptive tags to build a text prompt.
-For each training batch, we select between image and text conditioning modes by randomly dropping one modality.
-A drop probability of 75% for text and 25% for image empirically proved to be the most effective.
+StableMaterials allows flexible control over the generated material via text or image prompts describing high-level appearance.
+We encode the text or image condition using a pretrained CLIP model [Radford et al., 2021], which outputs a single feature vector.
+To make training robust, we alternate between image and text prompts.
+Specifically, we use (i) an ambient-lit rendering of the material as an image condition and (ii) a text caption for text prompts; when no caption is available, we generate short descriptive tags as an alternative.
+In each training batch, one modality is randomly dropped—text with 75% probability, image with 25%—to balance the two conditions.
 3.3. Semi-Supervised Adversarial Distillation
-To reduce the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model.
-However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties representation) during training.
-In particular, we complement the supervised loss with an unsupervised adversarial loss through a discriminator on the latent space ($LD$).
-This adversarial loss forces the generator to synthesize materials that present features similar to ground-truth ones, while the discriminator learns to distinguish between the two data sources.
-The generator is trained using samples from the material dataset as well as unannotated textures, while the discriminator is trained using only annotated samples.
-Our generator is optimized using a combination of supervised and adversarial losses. The supervised loss, $\mathcal{L}_{\text{sup}}$, is defined as:
-(1) $\mathcal{L}_{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\left[\left\|\epsilon_{t},\epsilon_{\theta}\left(z_{t,\text{mat}},t\right)\right\|^{2}\right]+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\left[\left\|\epsilon_{t},\epsilon_{\theta}\left(z_{t,\text{tex}},t\right)\right\|^{2}\right]$
-comparing the model-estimated noise $\epsilon\theta(z_{t},t)$ with the actual noise $\epsilon_{t}$ introduced during the forward diffusion process [Ho et al., 2020].
-Here, $z_{t,mat}$ and $z_{t,tex}$ are the noisy latent at timestep $t$ for the materials and textures samples respectively.
-The loss computed on unannotated texture samples is weighted using the hyperparameter $\alpha$, set to $0.15$ in our experiments.
-The adversarial objective $\mathcal{L}_{\text{adv}}$ push the generator to close the distribution gap between materials and textures, producing outputs that are perceived as materials by the discriminator.
-It is computed on the denoised latents at timestep $t-1$, and is defined as:
-(2) $\mathcal{L}_{\text{adv}}=-\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z})}\left[LD(z_{t-1})\right]$
-with $z_{t-1}=\text{concat}\left(z_{t-1,mat},z_{t-1,tex}\right)$ the concatenated denoised latents and $LD(\cdot)$ the output of the latent discriminator.
-Following Sauer et al.
-[2021, 2023a], $LD$ is trained using a hinge loss function [Lim and Ye, 2017] computed between the noised latent embeddings of the materials $z_{t-1,\text{real}}$ encoded with the VAE, and $z_{t-1,\text{fake}}$ the denoised latents from the generator.
-$LD$ shares the same architecture as the conditional diffusion U-Net encoder and is initialized with the weights of the diffusion U-Net.
-Functioning in a time-conditional manner, it incorporates the diffusion timestep and the CLIP embeddings from either text or image prompts, ensuring that evaluations are relevant to the desired material appearance.
 Figure 3.
@@
 Both the SDXL model and StableMaterials are prompted to generate the same material.
-The supervised $\mathcal{L}_{2}$ loss is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
+The supervised $\mathcal{L}_{2}$ loss, between the estimated noise and the added noise, is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
 […]
 Method Figure
+To bridge the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model.
+However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties) during training.
+Our method combines a supervised loss on annotated materials with an unsupervised adversarial loss that aligns the distribution of SDXL-generated textures with that of real materials in a shared latent space.
+The supervised term ensures correct reconstruction of ground-truth material properties, while the adversarial term guides the model to learn from unannotated textures by treating them as partially labeled data.
+To accomplish this, we introduce a latent discriminator ($LD$) that distinguishes between real material features and generated ones, effectively forcing the generator to produce material-like features even from unannotated textures.
+Our training combines:
+  1. (1)
+    Supervised Loss ($\mathcal{L}{\text{sup}}$): Ensures reconstruction of material properties, maintaining physical plausibility.
+  2. (2)
+    Adversarial Loss ($\mathcal{L}{\text{adv}}$): Guides the generator to map both materials and textures to a shared feature distribution, enabling diversity while maintaining realism.
+Our training strategy avoids mode collapse being primarely supervised, with the adversarial loss working as a distillation strategy.
+Supervised Loss.
+The supervised objective compares the denoiser’s prediction $\epsilon_{\theta}$ with the true noise $\epsilon_{t}$ introduced at each diffusion step $t$:
+(1) $\mathcal{L}{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon_{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{mat}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{tex}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}.$
+Here, $z_{t,\text{mat}}$ and $z_{t,\text{tex}}$ are noisy latents at step $t$ for materials and textures, respectively.
+The hyperparameter $\alpha$ (set to $0.15$) controls the relative importance of unannotated texture samples.
+Adversarial Loss.
+In parallel, $\mathcal{L}{\text{adv}}$ encourages the model to treat textures and materials similarly, guiding the generator to produce material-like outputs even when starting from unannotated textures.
+We compute this loss on the denoised latents $z_{t-1}$:
+(2) $\mathcal{L}{\text{adv}}=-\mathbb{E}{\mathbf{z}\sim p(\mathbf{z})}\Bigl{[}LD\bigl{(}z_{t-1}\bigr{)}\Bigr{]},$
+where $z_{t-1}=\text{concat}\Bigl{(}z_{t-1,\text{mat}},z_{t-1,\text{tex}}\Bigr{)}$ is the concatenation of denoised latents for materials and textures, and $LD(\cdot)$ is the output of our latent discriminator.
+By using the discriminator to align the distributions of material and texture latents, this loss effectively distills knowledge from SDXL-generated textures while ensuring they conform to the features of real materials.
+Latent Discriminator.
+We follow Sauer et al.
+[2021, 2023a] and train $LD$ with a hinge loss [Lim and Ye, 2017], comparing real latent embeddings $z_{t-1,\text{real}}$ (encoded from the VAE) against fake latent embeddings $z_{t-1,\text{fake}}$ (denoised by the generator).
+The discriminator operates in a time-conditional fashion: it receives the same timestep $t$ and the CLIP embedding as the generator.
+Architecturally, $LD$ mirrors the U-Net encoder of the diffusion model and is initialized with the same weights, leveraging its understanding of the latent space to effectively guide the generator.
+Unlike previous works that use adversarial distillation for fast generation [Sauer et al., 2023b, 2024], our approach bridges the domain gap between materials and textures.
+By training primarily with supervised terms while using adversarial guidance, the generator learns plausible material features even from unannotated textures.
+The adversarial component ensures SDXL-generated textures map to realistic material latents, effectively countering any shading artifacts and enriching the diversity of generated materials.
 3.4. Fast High-Resolution Generation
 Few steps generation
-To enhance generation efficiency, we finetune a Latent Consistency Model [Luo et al., 2023] (LCM).
-This model streamlines the denoising process by implementing a one-stage guided distillation of an augmented probability flow ordinary differential equation (PF-ODE).
-The distillation leverages a consistency function $f_{\theta}:(z_{t},c,t)\mapsto z_{0}$, which directly predicts the PF-ODE solution at $t=0$.
-Compared to two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the augmented PF-ODE.
-Additionally, LCM adopts a skipping step strategy, ensuring consistency not between consecutive steps, but between steps separated by $k$, thus $t_{n+k}\rightarrow t_{n}$, circumventing alignment issues in traditional models.
-The use of consistency models enables few-steps generation, resulting in fast material synthesis as detailed in Sec.
-4.2.
+To improve generation speed, we fine-tune a Latent Consistency Model (LCM) [Luo et al., 2023].
+LCM performs a one-stage guided distillation of an augmented Probability Flow ODE (PF-ODE) and directly predicts the solution at $t=0$ through a consistency function $f_{\theta}(z_{t},c,t)\mapsto z_{0}$.
+Unlike two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the PF-ODE and uses a skip-step strategy, ensuring consistency between time steps $t_{n+k}$ and $t_{n}$.
+This design avoids alignment issues present in typical two-stage approaches.
+As a result, LCM enables generation in only a few steps, resulting in faster material synthesis.
 Features rolling
-Recent material generation methods achieve tileability via noise rolling [Vecchio et al., 2023], leveraging the iterative nature of diffusion models.
-However, this approach becomes ineffective with fewer generation steps, being unable to fully remove seams.
-Circular padding [Zhou et al., 2022], on the other hand, despite being independent of diffusion steps, is not suitable for patched generation as it causes self-tileability in patches.
-To address these issues, we propose shifting tensor rolling from the diffusion step to the U-Net architecture by directly rolling the feature maps.
-This method, named features rolling, integrates rolling within U-Net blocks, reducing the need for multiple diffusion steps.
-By randomly translating and then reversing feature maps within each convolution and attention layer in the U-Net, this approach preserves edge continuity and minimizes visual artifacts typically associated with fewer diffusion steps.
-To better preserve structure in highly structured materials, we enable features rolling after the first diffusion step.
-We compare features rolling with other approaches to tileability in the Supplemental Materials.
+Noise rolling [Vecchio et al., 2024a] has been used to achieve tileability through iterative diffusion but becomes less effective with fewer steps.
+We address this limitation proposing the features rolling, which shifts rolling from the noisy inputs to the U-Net features.
+Within each convolution and attention layer, we randomly roll and then reverse the feature maps, thus preserving edge continuity while requiring fewer diffusion steps.
+For highly structured materials, we enable features rolling only after the first diffusion step to retain their global layout.
+We compare features rolling with other tileability methods in the Supplemental Materials.
 Latent Upscaling & Refinement
-Features rolling in combination with patch diffusion allows efficient high-resolution generation; however, it results in inconsistent generations between patches, due to the loss of cross-patch dependencies.
-To overcome this limitation, we use a two-stage pipeline, similar to the one introduced in SDXL [Podell et al., 2023], where the output of the base model is further refined by a high-resolution expert using SDEdit [Meng et al., 2021].
-This approach allows us to use patched diffusion in the refinement process while keeping the overall content consistent with the original generation.
-In particular, we first generate at the model’s base 512x512 resolution, and then refine it at the desired resolution.
-We apply the SDEdit refinement with a strength of $0.5$, which we find to be the best compromise between new high-frequency details cretion and consistency with the previous step generation.
-To better learn the high-frequency features of the materials, the refiner model is trained on 512x512 crops extracted from the 4K materials at their full resolution without downsampling.
-We show the quality improvement of this approach in the ablation study, Sec.
-4.4.3, and provide additional samples in Supplemental Materials.
+Combining features rolling with patch-based diffusion provides efficient, high-resolution synthesis, but patch-wise generation alone can introduce inconsistencies across tiles.
+We address this with a two-stage pipeline, similar to SDXL [Podell et al., 2023], using SDEdit [Meng et al., 2021] for refinement.
+Specifically, we first generate at $512\times 512$ resolution, then refine the output to the desired resolution with a strength of $0.5$, balancing new high-frequency details against global consistency.
+The refiner model is trained on full-resolution $512\times 512$ crops from 4K materials (no downsampling), ensuring it captures fine surface details.
+We demonstrate the effectiveness of this approach in Sec.
+4.4.3, with additional samples provided in the Supplemental Materials.
 4. Implementation & Results
@@
 [2021].
-We first train a multi-encoder VAE [Vecchio et al., 2024]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab.
-1).
+We first train a multi-encoder VAE [Vecchio et al., 2024b]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab.
+2).
 We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
@@
 Both models use the original OpenAI U-Net architecture with 18 input and output channels.
+Latent Consistency model
+The latent consistency model [Luo et al., 2023] is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$.
+We use a linear schedule for $\beta$ and denoise using the LCM (Latent Consistency Models) sampling schedule at inference time with $T=4$ steps.
+Inference
+We assess execution speed and memory usage.
+We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel at half-precision.
+Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM.
+In contrast, an LDM with a DDIM sampler (50 steps) plus 25 refinement steps requires 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
 Prompt Basecolor Normal Height Rough. Metallic Render Prompt Basecolor Normal Height Rough. Metallic Render
@@
 Full set of maps in Supplemental Materials.
-Latent Consistency model
-The latent consistency model [Luo et al., 2023] is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$.
-We use a linear schedule for $\beta$ and denoise using the LCM (Latent Consistency Models) sampling schedule at inference time with $T=4$ steps.
-Inference
-We assess execution speed and memory usage.
-We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel.
-Both base and refiner models run at half-precision.
-Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM.
-For comparison, the LDM with a DDIM sampler and 50 generation steps, followed by 25 refinement steps, takes 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
 4.3. Results and comparison
-For all results, we show generated material maps, as well as the rendering under ambient lighting.
-We include a wider variety results in the Supplemental Material, as well as material editing samples and a CLIP-based Nearest-Neighbour search in the training database.
+All results show both generated material maps and ambient-lit renderings.
+For evaluation, we carefully selected test prompts to include both in-domain concepts (present in training categories) and novel out-domain concepts, ensuring no direct overlap with training prompts.
+Additional samples, material editing results, and a CLIP-based nearest-neighbor search in our training database are included in the Supplemental Materials.
 Generation results
-We evaluate our model generation capabilities for both image (Fig.
-4 and Fig.
-12) and text (Fig.
-5 and Fig.
-13) conditioning.
-The results show the ability of StableMaterials to closely follow the input prompt, producing diverse realistic materials.
-For each set of results, we show two in-domain generations, that is materials categories that are present in our annotated training dataset, and two out-domain samples, materials from the unannotated data, confirming the ability of StableMaterials to generate new high-quality materials from both data distributions.
-Comparison
-We compare StableMaterials against MatFuse [Vecchio et al., 2024] trained on MatSynth [Vecchio and Deschaintre, 2024], MatGen [^1] Generation samples for MatGen were kindly provided by the authors.
-[Vecchio et al., 2023], Material Palette [Lopes et al., 2024], and the text-to-texture-to-material pipeline in Adobe Substance 3D Sampler [Adobe, 2024].
-Figures 6 and 7 show the generation results for image and text conditioning, respectively.
-It should be noted that MatFuse, Material Palette, and StableMaterials are trained on public datasets, while both MatGen and Sampler use private data.
-Furthermore, MatFuse, MatGen, and StableMaterials directly produce PBR maps, while MaterialPalette and Sampler employ a two-step process: first, generating a texture, then using an SVBRDF estimator to predict the material maps.
-This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exarcebated by the non-material-specific training sometimes generating natural images (with perspective) instead of surfaces.
+We show the generative capabilities of our model, for both image (Fig.
+4) and text (Fig.
+5) conditioning.
+We include both in-domain samples (categories found in the annotated dataset) and out-domain samples (from unannotated data).
+In all cases, StableMaterials produces realistic results that closely follow the prompts.
+Qualitative comparison
+We compare StableMaterials against MatFuse [Vecchio et al., 2024b], MatGen [Vecchio et al., 2024a], Material Palette [Lopes et al., 2024], and Adobe Substance 3D Sampler [Adobe, 2024], as shown in Figures 6 and 7 for image and text prompting, respectively.
+MatFuse, Material Palette, and StableMaterials each use public datasets, while MatGen and Sampler rely on private data.
+MatFuse is limited to a $256\times 256$ resolution, often generating blurry or simplistic outputs and struggling with complex textures.
+Material Palette and Sampler take a two-step approach (texture generation followed by SVBRDF estimation).This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exacerbated by the nonmaterial-specific training sometimes generating natural images (with perspective) instead of surfaces.
 Moreover, MaterialPalette fine-tunes a LoRA for each prompt, introducing a time and computation overhead.
-In contrast, StableMaterials directly creates material maps by having learned material features, leading to more precise and unbiased results.
-MatFuse is constrained by its $256\times 256$ resolution, resulting in blurry, undetailed generations, and struggles with complex materials.
-MatGen, on the other hand, generally synthesize high-quality materials, but presents oversharpening artifacts and struggles in following more complex promps.
+MatGen, on the other hand, produces high-quality materials, but presents over-sharpening artifacts and struggles in following more complex promps.
+In contrast, StableMaterials directly outputs PBR maps using a tailored latent representation and a two-stage generation process, mitigating resolution constraints and reducing artifacts.
 We only show the generation renderings in Figures 6 and 7, with the PBR maps and additional samples included in the Supplemental Materials.
+Quantitative Comparison
+Material quality is difficult to assess with standard image metrics (FID [Heusel et al., 2017], IS [Salimans et al., 2016]) due to the different data distribution from that of natural images.
+Instead, we leverage CLIP-based metrics (CLIP Score [Hessel et al., 2021] and CLIP-IQA [Wang et al., 2023]), which assess both semantic alignment with prompts and perceptual quality of outputs.
+Table 1 shows StableMaterials either outperforms or matches state-of-the-art methods—some trained on larger datasets—and significantly improves over MatFuse.
+Comparison is carried out on 80 text-conditioned generation.
+Table 1.
+Quantitative comparison.
+We compare the generation quality of StableMaterials to MatFuse, MatGen and Substance 3D Sampler.
+The CLIP-IQA score is computed using the “high-quality/low-quality” contrastive pair.
+MatFuse MatGen Substance 3D Sampler Stable Materials
+CLIP Score $\uparrow$ 26.2 28.8 24.9 29.6
+CLIP-IQA $\uparrow$ 0.52 0.66 0.71 0.70
 4.4. Ablation Study
@@
 We report the reconstruction performance, in terms of RMSE, for both the multi-encoder and single-encoder models in Tab.
-1.
+2.
 Results show that the adopted transfer-learning strategy allows us to obtain a smaller network while achieving performances comparable to those of the larger multiencoder VAE and retaining the same dientangled latent space.
 4.4.2. Training strategies
-We evaluate the semi-supervised training on model’s generation capabilities in Fig.
+We evaluate the effect of the semi-supervised training on model’s generation capabilities in Fig.
 8.
-The base model (w/o semi-supervision) generates high-quality in-domain material but, as expected, struggles with out-domain materials absent from the original training data.
-The semi-supervised approach, in contrast, enhances diversity and overall quality, allowing the generation of materials beyond the initial training scope.
+Our semi-supervised training improves generation diversity, especially on out-of-domain materials that the purely supervised baseline cannot produce reliably.
 In-domain samples Out-domain samples
-W/o Semi-supervision W/ Semi-supervision W/o Semi-supervision W/ Semi-supervision
+W/o distillation W/ distillation W/o distillation W/ distillation
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Semi-supervised learning improves generation quality and diversity, including new materials.
-Base Refined Base Refined
+Base Refined Patch diffusion Two-stages pipeline
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 9. Refinement ablation results. Generation crops: the use of a diffusion refiner significantly enhances generation quality and sharpness.
+(a) Refinement ablation.
+(b) High-resolution ablation.
+Figure 9.
+Ablation studies comparing (a) the effect of refinement on quality and (b) different approaches to achieving high-resolution.
+Results show that the use of a diffusion refiner significantly enhances generation quality and sharpness, and two-stage approach generates at the model’s native resolution avoids the inconsistency in scale of the patched diffusion.
 4.4.3. High-Resolution Generation
-We evaluate our method with and without the refiner in Fig.
-9.
-The generation looks sharper, less noisy, and more detailed.
-We demonstrate the effectiveness of our two-stage approach to high-resolution generation in Fig.
-10, comparing it to patched diffusion, complemented by feature rolling.
-Patched diffusion leads to increased scale (pattern replication) rather than enhanced resolution (more pixel density) and causes inconsistencies like scale variations between patches.
-In contrast, our approach generates consistent and better quality high-res materials by refining the generation at the model’s native resolution of $512\times 512$ pixels via a second diffusion step at the target resolution.
-Patch diffusion Two-stages pipeline
-Refer to caption Refer to caption
-‘Tortoise shell.’
-Figure 10.
-Ablation of the approaches to high-resolution.
-(1) Patched diffusion produces more scale rather than resolution and presents inconsistency in scale.
-(2) Our two-stage approach generates at the model’s native resolution, and refines the material.
+Our two-stage approach to high-resolution sharpens and improves details (Fig.9(a)).
+Additionally, using only patched diffusion can introduce scale and consistency artifacts across patches, whereas our two-stage approach (base generation + refinement) produces more coherent large-format outputs by consolidating patches at $512\times 512$ before upscaling (Fig.9(b)).
 5. Limitations and Future Work
 StableMaterials presents some limitations as shown in Fig.
-11.
-First, it struggles to handle natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures.
+10.
+First, it struggles with natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures.
 Introducing more variety in training prompts could help mitigate the problem.
@@
 The use of text prompts describing surface properties, at training time, could mitigate the problem.
+Finally, despite being able to represent a wide variety of materials outside the annotated dataset, it is still limited to the classes represented in the unannotated data.
 (1) (2) (3)
@@
 6. Conclusion
-We present StableMaterials, a novel diffusion model for fast generation of tileable high-resolution materials.
-Our model leverages semi-supervised training and knowledge distillation from large-scale pre-trained models to overcome the limitations in training data and improve material realism and variety.
-This approach significantly reduces reliance on annotated data, thereby enriching the diversity and realism of generated PBR materials.
-In conclusion, StableMaterials presents a robust solution to material generation challenges and we believe it can serve as a blueprint for future research to harness the potential of large-scale pretrained models and unsupervised data in computer graphics and beyond.
+We introduce StableMaterials, a novel diffusion-based model for fast, tileable, high-resolution material generation.
+By integrating semi-supervised training and knowledge distillation from large-scale pretrained models, StableMaterials overcomes the lack of annotated data and delivers greater realism and variety in PBR materials.
+Moreover, feature-rolling enables tileable generation in few-steps settings, making our approach practical for real-world applications.
+We believe that StableMaterials can serve as a blueprint for future research, demonstrating how domain-specific generators can effectively leverage large-scale pretrained models and unsupervised data to improve diversity while maintaining physical plausibility and generation quality.

### v4

@@
 StableMaterials: Enhancing Diversity in Material Generation via Semi-Supervised Learning
-Abstract.
+Abstract
 We introduce StableMaterials, a novel approach for generating photorealistic physical-based rendering (PBR) materials that integrate semi-supervised learning with Latent Diffusion Models (LDMs).
@@
 Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond.
-StableMaterials will be made publicly available.
-material appearance, generative models
-journal: TOGccs: Computing methodologies Appearance and texture representations
-Figure 1.
-We present StableMaterials, a diffusion–based model for materials generation through text or image prompting.
-Our approach enables high-resolution, tileable material maps, inferring both diffuse (Basecolor) and specular (Roughness, Metallic) properties, as well as the material mesostructure (Height, Normal).
-[…]
-Teaser Figure
-1. Introduction
+StableMaterials is publicly available at https://gvecchio.com/stablematerials.
+1 Introduction
 Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise.
-To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2024a, b].
+To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [6, 7, 31, 32, 39, 16, 65, 59, 3, 11, 60], or generation from a set of conditions [14, 17, 23, 66, 18, 60, 61].
 While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
 Despite recent efforts to create large-scale materials datasets, such as Deschaintre et al.
-[2018], OpenSVBRDF [Ma et al., 2023], and MatSynth [Vecchio and Deschaintre, 2024], these datasets are limited in diversity [Zhou et al., 2023], not capturing the vast range observed in large-scale image datasets such as LAION [Schuhmann et al., 2022].
+[6], OpenSVBRDF [38], and MatSynth [58], these datasets are limited in diversity [67], not capturing the vast range observed in large-scale image datasets such as LAION [55].
 These limitations can constrain the capabilities of learning-based approaches, potentially creating gaps in their generative capabilities and affecting realism and diversity.
 Fine-tuning has become a common practice to reduce these gaps in training data by leveraging existing knowledge from large-scale pretrained models.
-Techniques like Low-Rank Adaptation (LoRA) [Hu et al., 2021] effectively fine-tune models while preventing catastrophic forgetting.
-Methods such as Diff-Instruct [Luo et al., 2024], on the other hand, employ distillation strategies to transfer knowledge from pretrained models.
+Techniques like Low-Rank Adaptation (LoRA) [22] effectively fine-tune models while preventing catastrophic forgetting.
+Methods such as Diff-Instruct [37], on the other hand, employ distillation strategies to transfer knowledge from pretrained models.
 However, while fine-tuning or distillation within the same domain are straightforward, they pose significant challenges across different domains (e.g., image to material).
-To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [Podell et al., 2023] model.
+To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [44] model.
 In particular, we use a pretrained SDXL to generate unannotated texture samples from text prompts.
@@
 To include these textures in the training of our methods, we learn a common latent representation between textures and materials; then, we complement the traditional supervised loss, with an unsupervised adversarial loss, forcing the model to also generate realistic maps for unannotated samples and close the gap between the two data distributions.
-In addition, drawing inspiration from the SDXL [Podell et al., 2023] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation.
-We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [Meng et al., 2021] and patched diffusion.
+In addition, drawing inspiration from the SDXL [44] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation.
+We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [40] and patched diffusion.
 This approach allows to achieve high resolution while constraining the patched generation, ensuring consistency and memory efficiency.
-Subsequently, we distill a latent consistency model [Song et al., 2023] that allows fast generation by reducing the number of inference steps to four steps per stage.
-However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2024a] to achieve tileability.
+Subsequently, we distill a latent consistency model [57] that allows fast generation by reducing the number of inference steps to four steps per stage.
+However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [60] to achieve tileability.
 To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
@@
 The contributions of this work are as follows:
-  • StableMaterials, a new diffusion–based model for PBR material generation, leveraging a semi-supervised learning approach to incorporate unannotated data during training.
-  • A novel distillation technique to bridge the gap with large-scale models by including unannotated data in the training.
-  • A novel “features rolling” approach to tileability, minimizing the visual artifacts produced by fewer diffusion steps.
-  • State-of-the-art performance in PBR materials generation.
-2. Related Work
-Figure 2.
-Architecture of StableMaterials.
+  • StableMaterials, a new diffusion–based model for PBR material generation, leveraging semi-supervised learning to incorporate unannotated data during training.
+  • A novel adversarial distillation technique to bridge the gap with large-scale models.
+  • A novel “features rolling” approach to tileability, minimizing the artifacts produced by fewer diffusion steps.
+  • State-of-the-art performance in materials generation.
+2 Related Work
+Figure 1: Architecture of StableMaterials.
 The base model generates a low resolution materials of size 512x512.
-This generation is then upscaled and refined using SDEdit [Meng et al., 2021] by the refiner model using a patched approach to limit memory requirements.
-[…]
-Method Figure
+This generation is then upscaled and refined using SDEdit [40] by the refiner model using a patched approach to limit memory requirements.
 Materials Generation.
-Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a].
+Materials synthesis is an open challenge in computer graphics [13] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [6, 31, 32, 39, 16, 65, 59, 7, 3, 11, 60].
 Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al.
-[2020] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [Guo et al., 2020], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [Karras et al., 2020a]; Hu et al.
-[2022] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [Zhou et al., 2022], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training.
+[14] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [17], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [28]; Hu et al.
+[23] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [66], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training.
 Recent appraoches have focused on leveraging the generative capabilities of diffusion models for materials generation.
 In particular, Vecchio et al.
-[2024a] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image.
-MatFuse [Vecchio et al., 2024b] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting.
-MaterialPalette [Lopes et al., 2024] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [Hu et al., 2021] for each picture.
-Substance 3D Sampler [Adobe, 2024] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning.
+[60] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image.
+MatFuse [61] extends generation control with multimodal conditioning, and enables editing of existing materials via ‘volumetric’ inpainting.
+MaterialPalette [34] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [22] for each picture.
+Substance 3D Sampler [1] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning.
 However, these methods often lack diversity, struggling with complex material representation, or depend on image generation models, requiring additional steps to estimate the material parameters.
@@
 Generative models.
-Image generation is a long-standing challenge in computer vision, primarily due to the high dimensionality of images and complex data distributions.
-Generative Adversarial Networks (GAN) [Goodfellow et al., 2014] enabled the creation of high-quality images [Karras et al., 2017; Brock et al., 2018; Karras et al., 2020b], yet they suffer from unstable training [Arjovsky et al., 2017; Gulrajani et al., 2017; Mescheder, 2018], leading to mode collapse behaviour.
-Diffusion Models (DMs) [Sohl-Dickstein et al., 2015; Ho et al., 2020], particularly the more efficient Latent Diffusion (LDM) architecture [Rombach et al., 2022], have emerged as an alternative to GANs, achieving state-of-the-art results in image generation tasks [Dhariwal and Nichol, 2021], besides showing a more stable training.
-Following the success of LDMs, research has focused on improving the generation quality [Podell et al., 2023] and reducing the number of inference steps to speed up the generation process [Song et al., 2023; Luo et al., 2023; Sauer et al., 2023b, 2024] through model distillation.
-Furthermore, due to the proliferation of large-scale pretrained models, several approaches were proposed to reuse their knowledge [Hu et al., 2021; Luo et al., 2024; Ruiz et al., 2023].
+Image generation is a long-standing challenge in computer vision, due to the high dimensionality of images and complex data distributions.
+Generative Adversarial Networks (GAN) [12] enabled the creation of high-quality images [26, 4, 27], yet they suffer from unstable training [2, 15, 42] and mode collapse behaviour.
+Diffusion Models (DMs) [56, 21], particularly the more efficient Latent Diffusion (LDM) architecture [47], have emerged as an alternative to GANs, achieving state-of-the-art results in image generation tasks [8], besides showing a more stable training.
+Following the success of LDMs, research has focused on improving the generation quality [44] and reducing the number of inference steps to speed up the generation process [57, 36, 53, 54] through model distillation.
+Furthermore, due to the proliferation of large-scale pretrained models, several approaches were proposed to reuse their knowledge [22, 37, 49].
 However, these approaches focus on fine-tuning within the same image domain, restricting their applicability to non-image domains.
-3. Method
-StableMaterials builds on MatFuse [Vecchio et al., 2024b], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality, pixel-level reflectance properties for arbitrary materials (Fig.
-2).
-We replace MatFuse’s multi-encoder VAE architecture (used to learn a disentangled latent representation of material maps) with a more resource-efficient single-encoder model, fine-tuned to preserve the same latent properties.
-In addition, we introduce a semi-supervised training strategy that distills knowledge from a large-scale SDXL model to increase generation diversity.
-Our method leverages latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast, tileable generation.
+3 Method
+StableMaterials builds on MatFuse [61], which adapts the LDM paradigm [47] to synthesize high-quality, pixel-level reflectance properties for arbitrary materials (Fig.
+1).
+We replace MatFuse’s multi-encoder VAE architecture (used to learn a disentangled latent representation of material maps) with a more resource-efficient single-encoder model.
+We introduce a semi-supervised training strategy that distills knowledge from a large-scale SDXL model to increase generation diversity, and leverage latent consistency distillation [57] and a novel feature rolling technique for fast, tileable generation.
 A dedicated refiner model enables high-resolution outputs while preserving global consistency.
-3.1. Material Representation
-StableMaterials generates SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the mesostructure of the material.
+3.1 Material Representation
+StableMaterials generates SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [5, 25], using a GGX [62] distribution function, as well as the mesostructure of the material.
 Specifically, the generated maps include base color, normal, height, roughness, and metalness, where roughness specifies the specular lobe width and metalness indicates conductor regions.
-3.2. Material Generation
-The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
+3.2 Material Generation
+The generative model consists of a compression VAE [30] $\mathcal{E}$, encoding the material maps into a latent space, and a diffusion model [47] $\epsilon_{\theta}$, modeling the distribution of these latent features.
 Map Compression
@@
 Each encoder $\mathcal{E}_{i}$ encodes the $i^{th}$ map $\textbf{M}_{i}$ into a latent vector $z_{i}$, and the concatenated tensor $z=concat(z_{1},\dots,z_{N})$ is decoded to reconstruct the maps $\hat{M}=\mathcal{D}(z)$.
-Following [Vecchio et al., 2024b], training uses pixel-space $L_{2}$ loss, perceptual LPIPS [Zhang et al., 2021] loss, a patch-based adversarial loss [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018], with a Kullback-Leibler penalty [Kingma and Welling, 2013; Rezende et al., 2014] to regularize the latent space.
+Following [61], training uses pixel-space $L_{2}$ loss, perceptual LPIPS [64] loss, a patch-based adversarial loss [24, 9], and a rendering loss [6], with a Kullback-Leibler penalty [30, 46] to regularize the latent space.
 Afterward, we fine-tune a single-encoder model.
@@
 This preserves the disentangled latent representation and maintains compression efficiency with fewer parameters, as shown in Sec.
-4.4.1.
+4.4.
 Lastly, we create a shared latent space for texture and materials.
@@
 This model, based on Rombach et al.
-[2022], uses a time-conditional U-Net [Ronneberger et al., 2015] to denoise the latent vectors $z$.
+[47], uses a time-conditional U-Net [48] to denoise the latent vectors $z$.
 During training, we generate noised latent vectors using a deterministic forward diffusion process $q\left(z_{t}|z_{t-1}\right)$, as defined in Ho et al.
@@
 3.3
+Figure 2: Semi-Supervised training.
+Both the SDXL model and StableMaterials are prompted to generate the same material.
+The supervised $\mathcal{L}_{2}$ loss, between the estimated noise and the added noise, is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
 Conditioning
-StableMaterials allows flexible control over the generated material via text or image prompts describing high-level appearance.
-We encode the text or image condition using a pretrained CLIP model [Radford et al., 2021], which outputs a single feature vector.
+StableMaterials allows flexible generation control via text or image prompts describing high-level appearance.
+We encode the text or image condition using a pretrained CLIP model [45], which outputs a single feature vector.
 To make training robust, we alternate between image and text prompts.
@@
 In each training batch, one modality is randomly dropped—text with 75% probability, image with 25%—to balance the two conditions.
-3.3. Semi-Supervised Adversarial Distillation
-Figure 3.
-Semi-Supervised training.
-Both the SDXL model and StableMaterials are prompted to generate the same material.
-The supervised $\mathcal{L}_{2}$ loss, between the estimated noise and the added noise, is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
-[…]
-Method Figure
+3.3 Semi-Supervised Adversarial Distillation
 To bridge the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model.
 However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties) during training.
-Our method combines a supervised loss on annotated materials with an unsupervised adversarial loss that aligns the distribution of SDXL-generated textures with that of real materials in a shared latent space.
-The supervised term ensures correct reconstruction of ground-truth material properties, while the adversarial term guides the model to learn from unannotated textures by treating them as partially labeled data.
-To accomplish this, we introduce a latent discriminator ($LD$) that distinguishes between real material features and generated ones, effectively forcing the generator to produce material-like features even from unannotated textures.
-Our training combines:
-  1. (1)
-    Supervised Loss ($\mathcal{L}{\text{sup}}$): Ensures reconstruction of material properties, maintaining physical plausibility.
-  2. (2)
-    Adversarial Loss ($\mathcal{L}{\text{adv}}$): Guides the generator to map both materials and textures to a shared feature distribution, enabling diversity while maintaining realism.
-Our training strategy avoids mode collapse being primarely supervised, with the adversarial loss working as a distillation strategy.
-Supervised Loss.
+Our method combines a supervised loss on annotated materials with an unsupervised adversarial loss that aligns the latent distribution of SDXL-generated textures with that of real materials.
+The supervised Loss ($\mathcal{L}{\text{sup}}$) ensures correct reconstruction of ground-truth material properties, maintaining physical plausibility.
+The adversarial Loss ($\mathcal{L}{\text{adv}}$) guides the generator to map both materials and textures to a shared feature distribution, enabling diversity while maintaining realism.
+We introduce a latent discriminator ($LD$) that distinguishes between real material features and generated ones, effectively forcing the generator to produce material-like features even from unannotated textures.
+This strategy avoids mode collapse being primarely supervised, with the adversarial loss working as a distillation strategy.
+Supervised Loss
 The supervised objective compares the denoiser’s prediction $\epsilon_{\theta}$ with the true noise $\epsilon_{t}$ introduced at each diffusion step $t$:
-(1) $\mathcal{L}{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon_{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{mat}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{tex}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}.$
+$\mathcal{L}_{\text{sup}}$ $=\mathbb{E}_{t,z_{0},\epsilon}\left[\lVert\epsilon_{t}-\epsilon_{\theta}(z_{t,\text{mat}},t)\rVert^{2}\right]$ (1)
+$\quad+\alpha\,\mathbb{E}_{t,z_{0},\epsilon}\left[\lVert\epsilon_{t}-\epsilon_{\theta}(z_{t,\text{tex}},t)\rVert^{2}\right].$ (2)
 Here, $z_{t,\text{mat}}$ and $z_{t,\text{tex}}$ are noisy latents at step $t$ for materials and textures, respectively.
 The hyperparameter $\alpha$ (set to $0.15$) controls the relative importance of unannotated texture samples.
-Adversarial Loss.
+Adversarial Loss
 In parallel, $\mathcal{L}{\text{adv}}$ encourages the model to treat textures and materials similarly, guiding the generator to produce material-like outputs even when starting from unannotated textures.
 We compute this loss on the denoised latents $z_{t-1}$:
-(2) $\mathcal{L}{\text{adv}}=-\mathbb{E}{\mathbf{z}\sim p(\mathbf{z})}\Bigl{[}LD\bigl{(}z_{t-1}\bigr{)}\Bigr{]},$
+$\mathcal{L}{\text{adv}}=-\mathbb{E}{\mathbf{z}\sim p(\mathbf{z})}\Bigl[LD\bigl(z_{t-1}\bigr)\Bigr],$ (3)
 where $z_{t-1}=\text{concat}\Bigl(z_{t-1,\text{mat}},z_{t-1,\text{tex}}\Bigr)$ is the concatenation of denoised latents for materials and textures, and $LD(\cdot)$ is the output of our latent discriminator.
 By using the discriminator to align the distributions of material and texture latents, this loss effectively distills knowledge from SDXL-generated textures while ensuring they conform to the features of real materials.
-Latent Discriminator.
+Latent Discriminator
 We follow Sauer et al.
-[2021, 2023a] and train $LD$ with a hinge loss [Lim and Ye, 2017], comparing real latent embeddings $z_{t-1,\text{real}}$ (encoded from the VAE) against fake latent embeddings $z_{t-1,\text{fake}}$ (denoised by the generator).
+[51, 52] and train $LD$ with a hinge loss [33], comparing real latent embeddings $z_{t-1,\text{real}}$ (encoded from the VAE) against fake latent embeddings $z_{t-1,\text{fake}}$ (denoised by the generator).
 The discriminator operates in a time-conditional fashion: it receives the same timestep $t$ and the CLIP embedding as the generator.
 Architecturally, $LD$ mirrors the U-Net encoder of the diffusion model and is initialized with the same weights, leveraging its understanding of the latent space to effectively guide the generator.
-Unlike previous works that use adversarial distillation for fast generation [Sauer et al., 2023b, 2024], our approach bridges the domain gap between materials and textures.
+Only ground-truth material latents are treated as real samples by the discriminator.
+Unlike previous works that use adversarial distillation for fast generation [53, 54], our approach bridges the domain gap between materials and textures.
 By training primarily with supervised terms while using adversarial guidance, the generator learns plausible material features even from unannotated textures.
 The adversarial component ensures SDXL-generated textures map to realistic material latents, effectively countering any shading artifacts and enriching the diversity of generated materials.
-3.4. Fast High-Resolution Generation
+3.4 Fast High-Resolution Generation
 Few steps generation
-To improve generation speed, we fine-tune a Latent Consistency Model (LCM) [Luo et al., 2023].
+To improve generation speed, we fine-tune a Latent Consistency Model (LCM) [36].
 LCM performs a one-stage guided distillation of an augmented Probability Flow ODE (PF-ODE) and directly predicts the solution at $t=0$ through a consistency function $f_{\theta}(z_{t},c,t)\mapsto z_{0}$.
-Unlike two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the PF-ODE and uses a skip-step strategy, ensuring consistency between time steps $t_{n+k}$ and $t_{n}$.
+Unlike two-stage methods [41], LCM integrates the guidance scale $\omega$ directly into the PF-ODE and uses a skip-step strategy, ensuring consistency between time steps $t_{n+k}$ and $t_{n}$.
 This design avoids alignment issues present in typical two-stage approaches.
@@
 Features rolling
-Noise rolling [Vecchio et al., 2024a] has been used to achieve tileability through iterative diffusion but becomes less effective with fewer steps.
+Noise rolling [60] has been used to achieve tileability through iterative diffusion but becomes less effective with fewer steps.
 We address this limitation proposing the features rolling, which shifts rolling from the noisy inputs to the U-Net features.
@@
 For highly structured materials, we enable features rolling only after the first diffusion step to retain their global layout.
-We compare features rolling with other tileability methods in the Supplemental Materials.
+We compare features rolling with other tileability methods in Supplemental Materials.
 Latent Upscaling & Refinement
 Combining features rolling with patch-based diffusion provides efficient, high-resolution synthesis, but patch-wise generation alone can introduce inconsistencies across tiles.
-We address this with a two-stage pipeline, similar to SDXL [Podell et al., 2023], using SDEdit [Meng et al., 2021] for refinement.
+We address this with a two-stage pipeline, similar to SDXL [44], using SDEdit [40] for refinement.
 Specifically, we first generate at $512\times 512$ resolution, then refine the output to the desired resolution with a strength of $0.5$, balancing new high-frequency details against global consistency.
@@
 We demonstrate the effectiveness of this approach in Sec.
-4.4.3, with additional samples provided in the Supplemental Materials.
-4. Implementation & Results
+4.4, with additional samples provided in Supplemental Materials.
+4 Implementation & Results
 In this section, we first introduce the datasets used in our work, we then assess the generation capabilities of StableMaterials and compare it with recent state-of-the-art methods.
-Finally, we evaluate our design choices in the ablation studies.
-4.1. Dataset
-We train our model on the combined MatSynth [Vecchio and Deschaintre, 2024] and Deschaintre et al.
+We evaluate our design choices in the ablation studies.
+4.1 Dataset
+We train our model on the combined MatSynth [58] and Deschaintre et al.
 [6] datasets, for a total of $6,198$ unique PBR materials, using the original training/test splits.
 Our dataset includes 5 material maps (Basecolor, Normal, Height, Roughness, Metallic) and their renderings under different environmental illuminations.
-We complement the dataset with $4,000$ texture-text pairs from SDXL [Podell et al., 2023], using $200$ prompts.
-We query a ChatGPT [OpenAI, 2024] model to suggest relevant material prompts.
-The full list of prompts used, as well as samples generated from SDXL, are provided in the Supplemental Material.
-4.2. Technical details
+We complement the dataset with $4,000$ texture-text pairs from SDXL [44], using $200$ prompts.
+We query a ChatGPT [43] model to suggest relevant material prompts.
+The full list of prompts and the SDXL textures, are provided in Supplemental Material.
+4.2 Technical details
 We train all models on a single NVIDIA RTX4090 GPU with 24GB of VRAM, employing gradient accumulation to achieve a larger batch.
-Autoencoder
-The compression model is trained with mini-batch gradient descent, using the Adam optimizer [Kingma and Ba, 2014] and a batch size of $8$.
+Autoencoder is trained with mini-batch gradient descent, using the Adam optimizer [29] and a batch size of $8$.
 We train the model for $1,000,000$ iterations with a learning rate of $4.5\cdot 10^{-4}$ and enable the $\mathcal{L}_{\text{adv}}$ after $300,000$ iterations as in Esser et al.
 [10].
-We first train a multi-encoder VAE [Vecchio et al., 2024b]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab.
+We first train a multi-encoder VAE [61]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab.
 2).
 We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
-Latent Diffusion model
-The diffusion model is trained supervisedly for $400,000$ iterations with a batch size of $16$ using an AdamW [Loshchilov and Hutter, 2017] optimizer, with a learning rate of $3.2\cdot 10^{-5}$.
-The model is fine-tuned semi-supervisedly for $200,000$ iterations with a batch size of $8$.
-We fine-tune the refiner model for $50,000$ iterations using a batch size of $16$ and a learning rate of $2\cdot 10^{-7}$.
-Both models use the original OpenAI U-Net architecture with 18 input and output channels.
-Latent Consistency model
-The latent consistency model [Luo et al., 2023] is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$.
-We use a linear schedule for $\beta$ and denoise using the LCM (Latent Consistency Models) sampling schedule at inference time with $T=4$ steps.
-Inference
-We assess execution speed and memory usage.
-We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel at half-precision.
-Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM.
-In contrast, an LDM with a DDIM sampler (50 steps) plus 25 refinement steps requires 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
 Prompt Basecolor Normal Height Rough. Metallic Render Prompt Basecolor Normal Height Rough. Metallic Render
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 4.
-Image-prompting.
+Figure 3: Image-prompting.
 StableMaterials accurately captures the visual appearance of the input image, producing realistic materials for both in-domain (on the left) and out-domain (on the right) prompts.
@@
 ‘Old wooden parquet floor.’ ‘Batik fabric with Indonesian patterns.’
-Figure 5.
-Text-prompting.
+Figure 4: Text-prompting.
 StableMaterials closely follow the input prompt, producing realistic materials for both in-domain (on the left) and out-domain (on the right) samples.
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 6.
-Comparison for image-prompting.
+Figure 5: Comparison for image-prompting.
 We compare StableMaterials with MatFuse, MatGen, and MaterialPalette in image-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model.
 StableMaterials improves over previous methods quality and ability to captures the visual appearance of the input image.
-Full set of maps in Supplemental Materials.
 MatFuse MatGen Substance 3D Sampler Stable Materials MatFuse MatGen Substance 3D Sampler Stable Materials
@@
 ‘Volcanic lava.’ ‘Coral pavona with a bumpy and porous texture.’
-Figure 7.
-Comparison for text prompting.
+Figure 6: Comparison for text prompting.
 We compare StableMaterials with MatFuse, MatGen, and Substance 3D Sampler on text-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model.
@@
 Visual quality is on par or better than models trained on larger datasets.
-Full set of maps in Supplemental Materials.
-4.3. Results and comparison
+Latent Diffusion model is trained supervisedly for $400,000$ iterations with a batch size of $16$ using an AdamW [35] optimizer, with a learning rate of $3.2\cdot 10^{-5}$.
+The model is fine-tuned semi-supervisedly for $200,000$ iterations with a batch size of $8$.
+We fine-tune the refiner model for $50,000$ iterations using a batch size of $16$ and a learning rate of $2\cdot 10^{-7}$.
+Both models use the original OpenAI U-Net architecture with 18 input and output channels.
+Latent Consistency model is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$.
+We use a linear schedule for $\beta$ and denoise using the LCM sampling schedule at inference time with $T=4$ steps.
+Inference
+We assess execution speed and memory usage.
+We generate with 4 denoising steps, followed by 2 refinement steps, using the LCM sampler with a fixed seed and processing up to 8 patches in parallel at half-precision.
+Generation takes 0.6s at $512\times 512$, 1.5s at $1024\times 1024$ and 6.5GB of VRAM, 4.9s at $2048\times 2048$ and 7.4GB VRAM, and 18.6s at $4096\times 4096$ and 12GB of VRAM.
+In contrast, an LDM with a DDIM sampler (50 steps) plus 25 refinement steps requires 20.6s at $2048\times 2048$, and 65.4s at $4096\times 4096$.
+4.3 Results and comparison
 All results show both generated material maps and ambient-lit renderings.
 For evaluation, we carefully selected test prompts to include both in-domain concepts (present in training categories) and novel out-domain concepts, ensuring no direct overlap with training prompts.
-Additional samples, material editing results, and a CLIP-based nearest-neighbor search in our training database are included in the Supplemental Materials.
+Additional samples, material editing results, and a CLIP-based nearest-neighbor search are included in the Supplemental Materials.
 Generation results
 We show the generative capabilities of our model, for both image (Fig.
-4) and text (Fig.
-5) conditioning.
+3) and text (Fig.
+4) conditioning.
 We include both in-domain samples (categories found in the annotated dataset) and out-domain samples (from unannotated data).
@@
 Qualitative comparison
-We compare StableMaterials against MatFuse [Vecchio et al., 2024b], MatGen [Vecchio et al., 2024a], Material Palette [Lopes et al., 2024], and Adobe Substance 3D Sampler [Adobe, 2024], as shown in Figures 6 and 7 for image and text prompting, respectively.
+We compare StableMaterials against MatFuse [61], MatGen [60], Material Palette [34], and Adobe Substance 3D Sampler [1], as shown in Figures 5 and 6 for image and text prompting, respectively.
 MatFuse, Material Palette, and StableMaterials each use public datasets, while MatGen and Sampler rely on private data.
@@
 In contrast, StableMaterials directly outputs PBR maps using a tailored latent representation and a two-stage generation process, mitigating resolution constraints and reducing artifacts.
-We only show the generation renderings in Figures 6 and 7, with the PBR maps and additional samples included in the Supplemental Materials.
+We only show the generation renderings in Figures 5 and 6, with the PBR maps and additional samples included in the Supplemental Materials.
+User Study
+To validate the quality of our results, we conducted a user study with 20 domain experts.
+Participants compared material generations from the different methods across ten prompts.
+For each prompt, they selected their preferred sample and rated all four materials on a 1–5 scale for prompt fidelity, realism, and material consistency.
+StableMaterials was preferred 105 times, compared to 66 for Sampler, 25 for MatGen, and 4 for MatFuse.
+It also achieved the highest average rating (3.50), outperforming Sampler (2.71), MatGen (2.28), and MatFuse (1.85).
+A one-way ANOVA confirmed the differences are statistically significant ($F=60.5$, $p<0.05$).
+A chi-square goodness-of-fit test on the preference counts $[4,25,66,105]$ against a uniform distribution yielded $\chi^{2}=120.44$, $p<0.001$, indicating strong deviation from chance.
+A correlation coefficient of $R=0.99$ confirms alignment between average ratings and preference counts.
 Quantitative Comparison
-Material quality is difficult to assess with standard image metrics (FID [Heusel et al., 2017], IS [Salimans et al., 2016]) due to the different data distribution from that of natural images.
-Instead, we leverage CLIP-based metrics (CLIP Score [Hessel et al., 2021] and CLIP-IQA [Wang et al., 2023]), which assess both semantic alignment with prompts and perceptual quality of outputs.
+Material quality is difficult to assess with standard image metrics (FID [20], IS [50]) due to the different data distribution from that of natural images.
+Instead, we leverage CLIP-based metrics (CLIP Score [19] and CLIP-IQA [63]), which assess both semantic alignment with prompts and perceptual quality of outputs.
 Table 1 shows StableMaterials either outperforms or matches state-of-the-art methods—some trained on larger datasets—and significantly improves over MatFuse.
 Comparison is carried out on 80 text-conditioned generation.
-Table 1.
-Quantitative comparison.
-We compare the generation quality of StableMaterials to MatFuse, MatGen and Substance 3D Sampler.
-The CLIP-IQA score is computed using the “high-quality/low-quality” contrastive pair.
 MatFuse MatGen Substance 3D Sampler Stable Materials
@@
 CLIP-IQA $\uparrow$ 0.52 0.66 0.71 0.70
-4.4. Ablation Study
+Table 1: Quantitative comparison.
+We compare the generation quality of StableMaterials to MatFuse, MatGen and Substance 3D Sampler.
+The CLIP-IQA score is computed using the “high-quality/low-quality” contrastive pair.
+4.4 Ablation Study
 We evaluate our different design choices by comparing the performance of our model against the baseline solutions. We provide high-resolution ablation results in the Supplemental Materials.
-4.4.1. VAE Architecture
+VAE Architecture
+Albedo Normal Height Rough. Metal.
+Multi-$\mathcal{E}$ (271M par.) 0.030 0.035 0.030 0.032 0.016
+Single-$\mathcal{E}$ (101M par.) 0.028 0.037 0.030 0.032 0.015
 […]
-Basecolor Normal Height Rough. Metallic
-$-\mathcal{E}$ (271M par.) 0.030 0.035 0.030 0.032 0.016
-$-\mathcal{E}$ (101M par.) 0.028 0.037 0.030 0.032 0.015
 We report the reconstruction performance, in terms of RMSE, for both the multi-encoder and single-encoder models in Tab.
@@
 Results show that the adopted transfer-learning strategy allows us to obtain a smaller network while achieving performances comparable to those of the larger multiencoder VAE and retaining the same dientangled latent space.
-4.4.2. Training strategies
+Training strategies
 We evaluate the effect of the semi-supervised training on model’s generation capabilities in Fig.
-8.
+7.
 Our semi-supervised training improves generation diversity, especially on out-of-domain materials that the purely supervised baseline cannot produce reliably.
+High-Resolution Generation
+Our two-stage approach to high-resolution sharpens and improves details (Fig.8(a)).
+Additionally, using only patched diffusion can introduce scale and consistency artifacts across patches, whereas our two-stage approach (base generation + refinement) produces more coherent large-format outputs by consolidating patches at $512\times 512$ before upscaling (Fig.8(b)).
+5 Limitations and Future Work
+StableMaterials presents some limitations as shown in Fig.
+9.
+First, it struggles with natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures.
+Introducing more variety in training prompts could help mitigate the problem.
+Additionally, it can occasionally generate incorrect reflectance properties (e.g.: material misclassified as metal) for material classes that are only present in the unannotated dataset.
+The use of text prompts describing surface properties, at training time, could mitigate the problem.
+Finally, despite being able to represent a wide variety of materials outside the annotated dataset, it is still limited to the classes represented in the unannotated data.
 In-domain samples Out-domain samples
@@
 ‘Terracotta brick wall with white grout.’ ‘Kilim Rugs carpets with complex geometric patterns.’
-Figure 8.
-Ablation study of the training stategies.
+Figure 7: Ablation study of the training stategies.
 The model effectively generates high-quality material with image or text prompts but struggles with unrepresented materials.
@@
 (b) High-resolution ablation.
-Figure 9.
-Ablation studies comparing (a) the effect of refinement on quality and (b) different approaches to achieving high-resolution.
+Figure 8: Ablation studies comparing (a) the effect of refinement on quality and (b) different approaches to achieving high-resolution.
 Results show that the use of a diffusion refiner significantly enhances generation quality and sharpness, and two-stage approach generates at the model’s native resolution avoids the inconsistency in scale of the patched diffusion.
-4.4.3. High-Resolution Generation
-Our two-stage approach to high-resolution sharpens and improves details (Fig.9(a)).
-Additionally, using only patched diffusion can introduce scale and consistency artifacts across patches, whereas our two-stage approach (base generation + refinement) produces more coherent large-format outputs by consolidating patches at $512\times 512$ before upscaling (Fig.9(b)).
-5. Limitations and Future Work
-StableMaterials presents some limitations as shown in Fig.
-10.
-First, it struggles with natural prompts describing spatial relations and is unable to accurately represent complex concepts or figures.
-Introducing more variety in training prompts could help mitigate the problem.
-Additionally, it can occasionally generate incorrect reflectance properties (e.g.: material misclassified as metal) for material classes that are only present in the unannotated dataset.
-The use of text prompts describing surface properties, at training time, could mitigate the problem.
-Finally, despite being able to represent a wide variety of materials outside the annotated dataset, it is still limited to the classes represented in the unannotated data.
 (1) (2) (3)
@@
 ‘Squared tiles enclosed by rectangular tiles.’ ‘Yukata fabric with patterns of dragons.’ ‘Shoji Screens made of paper and wood.’
-Figure 10. Limitations. Left to right: (1) Struggles with complex prompts describing spatial relations. (2) Unable to represent complex figures or patterns. (3) Can hallucinate reflectance properties.
-6. Conclusion
+Figure 9: Limitations. Left to right: (1) Struggles with complex prompts describing spatial relations. (2) Unable to represent complex figures or patterns. (3) Can hallucinate reflectance properties.
+6 Conclusion
 We introduce StableMaterials, a novel diffusion-based model for fast, tileable, high-resolution material generation.
-By integrating semi-supervised training and knowledge distillation from large-scale pretrained models, StableMaterials overcomes the lack of annotated data and delivers greater realism and variety in PBR materials.
-Moreover, feature-rolling enables tileable generation in few-steps settings, making our approach practical for real-world applications.
+By integrating semi-supervised training and knowledge distillation from large-scale pretrained models, it overcomes the lack of annotated data and delivers greater realism and variety in PBR materials.
+Moreover, features rolling enables tileable generation in few-steps settings, making our approach practical for real-world applications.
 We believe that StableMaterials can serve as a blueprint for future research, demonstrating how domain-specific generators can effectively leverage large-scale pretrained models and unsupervised data to improve diversity while maintaining physical plausibility and generation quality.